Operations Report – April 16th

Crew 192 Operations Report 16Apr2018

SOL 8

Name of person filing report: Richard Blakeman

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 78%

Diesel – 70% fuel

Propane – 63% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Approximately 4 gallons left in one and the other is empty

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 780 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133570.4

Toilet tank emptied: Yes. Emptied today, April 16.

ATV’s Used: None

Oil Added? No but checked and OK.

ATV Fuel Used: 5%

# Hours the ATVs were Used today: 0.5

Notes on ATVs: All working well and ready for use, gas filled up yesterday.

Deimos rover used: Yes
Hours: 120
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 38.2
Beginning charge: 100%
Ending charge: 98%
Currently charging: Yes

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.9
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well, though the bad weather is hampering our EVAs and astronomical activities.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, 3 two-piece suits used, radios working fine.

Summary of Hab operations: Finished removing hab skirt as requested, kept a close eye on all hab systems as strong winds blew most of the day.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water remains in the RAM for continued cleaning of embedded sediment.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Requesting trash pickup.

EVA Report – April 16th

EVA Report

EVA #10 – 16Apr2018 – Sol 8

Crew members: Joe Dituri, Abdreea Radulescu, Victoria Varone

Objectives:
To hike along the ridge and see the overview of the hab as well as compare fogging of suits with heavy exertion.

Operations:
0730-1030

Narrative: Three crew members took a hike along walking path 1103. This was hard work and fogged the suits badly. We walked up the ridge and then proceeded to the far ridge to the west but we did not achieve it .

Vehicles
None

__________

EVA #11 16 April 2018 – Sol 8
Crew Members: Richard Blakeman and Ashok Narayanamoorthi.

Objectives: Originally planned to explore south Lowell road and near route 1101. Scrubbed mission for safety and high winds.

Operations: 1500-1600

Narrative: Took Diemos Rover and put both crew in the rover to deliver the mouse to a distance far enough away so that he would not come back…

Vehicles
Diemos Rover

Greenhab Report – April 16th

GreenHab Report
Name: Andreea Radulescu
Date: 16 Apr 2018

Environmental control:

Ambient with door open
Shade cloth: on
Fan: not on.

Average temperatures: 27.1℃
Low 13.7℃
High 53.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 8 gal

Time(s) of watering for crops: Twice: In the morning around 10:30 am and afternoon 4:45pm.

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: No significant changes. Picked another batch of tomatoes for the crew.

Support/supplies needed: None

Sol Summary – April 16th

Crew 192 Sol Summary Report 16Apr2018

Sol 8

Summary Title: Houston, We Have a [Mouse] Problem

Author: Victoria Varone

Mission Status:

Sol Activity Summary: Our day started with a visit from a tiny companion ‒ code name “Paul Watney” in honor of our missing crew member and “The Martian” ‒ who made his way into the upper deck of the hab. He accompanied us in the hab for the day but then decided to return to Mars. (see Photo Report)

Our first EVA of the day was a long walking hike along the ridge. We climbed some of the dunes and hills and had to get creative along the way as there weren’t outlets or shallow spots where we thought they would be. One of the highlights of the hike was making our way up to the hab overview, which offered some amazing Martian views overlooking the hab.

Our second EVA was shortened due to strong winds that increased as the day went one. It was turned into a short relocation EVA where we dropped off our tiny friend “Paul Watney” back onto the vast plains of Mars.

The evening consisted of some more cooking for the crew, baking peanut butter cookies, and watching “Apollo 13.”

Look Ahead Plan: Finish cleaning Water Tank and final skirt HAB clean up. Exploration continues to the south and spreading North.

Anomalies in Work: None

Weather: Chilly early on, warmer later in the day, mostly cloudy all day.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report, GreenHab Report

Support Requested: None

Maintenance request: See Operations Report

Commander Report – April 16th

Joseph Dituri

16 April 18

TITLE: There was a mouse about the house.

Last night the crew decided to play Coupe. This card game had us rolling for several hours. “Duke Money” was yelled myriad times and we fondly recalled our other team mates.

When we went to sleep one person noted a mouse going under the frig. XO set out a trap and we had a visitor by morning. We named him Paul Watney for our missing crew member and the Martian.

The morning EVA was a walking expedition to test exertion and fogging. We also hiked the ridge behind the HAB and made it to almost the next ridge…death march by Andreea.

Upon return we had a super brunch and the crew who stayed field day-ed the upper portion of the HAB …slowley we are moving every piece of furnature and cleaning behind or under it so that this place will be super clean at the end.

Afternoon EVA was curtailed….we changed it to a short EVA mission to drop off mouse Watney far enough away that he could not walk back. I did an Operational Risk Mitigation matrix and decided it was safe to proceed. We put two people in one Rover and proceeded. The wind kicked up a bit but it was safe and we got rid of the mouse. RTB coupled with some minor engineering tasks to ensure the cargo on Mars Society trailer would not fly free from increased wind.

Uneventful dinner and rest of the day …planning on resting a bit. We made peanut butter cookies. Yum.

IMPRESSED: I was impressed with Victoria ( Blue 04) who persevered on a death march up hill BOTH WAYS and got the best shit of the trip so far….what a photographer.

LEARNED: I leanred that the map at our disposal is more or less a “guide”….if it says “space suit road” that translates to unrecognizable path that goes up the hill…;*)

POA&M: Tomorrow keep cleaning the water tank and finish HAB skirt removal process. EVAs to look north and finish up south exploration.

~Dituri sends